# envname-P envname#1

### v1

@@ 1 Introduction
As Large Language Models (LLMs) grow in size and on-device applications gain traction (Wahab & Adda, 2025), significant attention has been devoted to reducing inference costs via model compression (Frantar et al., 2022; Lin et al., 2024; Ma et al., 2023). One state-of-the-art method is quantization-aware training (QAT) (Chen et al., 2025a; Lin et al., 2024; Liu et al., 2025; Jacob et al., 2018). To adapt the model to the loss of numerical precision, QAT incorporates quantization directly into the model training process. It has been shown that QAT outperforms post-training quantization (PTQ) (Xiao et al., 2023; Banner et al., 2019), where quantization is applied after training is completed. Moreover, Liu et al. (2025) demonstrated that the best accuracy is achieved when a QAT phase follows a full-precision (FP) training phase.
@@
  • We propose a novel approach: cooldown & QAT fusion—a scheme where learning rate decay is performed jointly with quantization-aware training, eliminating redundant full-precision updates and achieving better accuracy for the same token count.
@@ 2 Related Work
Multiple works have previously addressed the problem of predicting final model loss ($L$) as a function of parameter count ($N$) and consumed tokens ($D$) (Bi et al., 2024; Hoffmann et al., 2022b; Kaplan et al., 2020). The Chinchilla (Hoffmann et al., 2022b) loss model is one of the most commonly used: $L(N,D)=E+AN^{-\alpha}+CD^{-\beta},$ where $A$, $C$, $\alpha$, $\beta$, and $E$ are fitted parameters. Bi et al. (2024) expand on this idea, fitting accuracy as a function of used non-embedding FLOPs (FLOP estimation of model inference without embedding layer calculations), showing that such an approach works better across different model sizes. Additionally, they show that scaling laws are greatly influenced by data quality.
@@
3 Experimental Setup
We use a decoder-only transformer (Zhang et al., 2021) identical to Llama 2 (Touvron et al., 2023). The architecture incorporates SwiGLU activations (Shazeer, 2020), RoPE (Su et al., 2024), RMSNorm (Zhang & Sennrich, 2019), alternating attention and feed-forward layers, and tied embedding and language-modeling head weights. We use the Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.99,\varepsilon=10^{-8}$) with decoupled weight decay of $0.01$ (Loshchilov & Hutter, 2017; Dereich et al., 2024) for all parameters outside the embedding and normalization layers. All experiments are trained with bfloat16 automatic mixed precision (Liu et al., 2021). Training is conducted on the DCLM dataset (Li et al., 2024), tokenized with the Llama 2 tokenizer with a 32,000-token vocabulary. We merge all tokenized documents into a single sequence with appropriate delimiting tokens and take chunks of 1024 tokens (used sequence length) for the batch—an approach also known as “concat-and-chunk” (Pouransari et al., 2024). The dataset is split into training and validation sets, and validation perplexity is used for evaluation. For QAT algorithms, we rely on ParetoQ (Liu et al., 2025) for our setups, as this method achieves state-of-the-art accuracy across different bit widths by combining different approaches. Additional information on the training procedure is provided in appendix 9.
4 Optimal QAT Compute
To study how loss changes for different combinations of QAT/FP training length, we train models of different sizes, different FP stage token counts ($D_{\text{fp}}$), QAT token counts ($D_{\text{qat}}$), and different QAT bit widths $B$. For the smallest model (86.0M parameters), we conduct experiments from 2.3B to 1.4T total tokens, while for the largest model (759.0M parameters), we conduct experiments from 8.5B to 669.2B total tokens. We focus primarily on 1-, 2-, 4-, and 6-bit quantization widths. Token counts and QAT fractions for our experiments are reported in appendix 21. We also verify that the obtained post-QAT models maintain reasonable accuracy and present a comparison to full-precision models in appendix 22. Specifically, our 4- and 6-bit setups achieve quality close to that of the full-precision model for the same total token count, and the drop in quality for 1- and 2-bit is reasonable.
@@ 4 Optimal QAT Compute
4.1 Predicting the Optimal QAT Fraction
In this section, we focus on fitting the optimal QAT fractions directly. To account for different QAT bit widths used, we introduce the tokens-per-parameter-byte statistic. This choice was made based on several observations: larger models are generally easier to quantize, models trained for longer are harder to quantize, and smaller QAT bit widths are harder to quantize as well. While being intuitive from a QAT accuracy perspective, it can also predict the optimal fraction with high precision. Figure 2 provides a comparison between the two approaches. It is clearly seen that using tokens-per-parameter-byte provides an interpretable adjustment, facilitating a better fit. For optimal QAT fraction prediction, we fit a function of the form
@@ 4.1 Predicting the Optimal QAT Fraction
where $\widehat{f}$ is the predicted optimal QAT fraction for the total tokens-per-parameter-byte count $S_{\text{total}}$, and $a$ is a fitted parameter. This function choice was made due to the observed almost linear dependency in log-log coordinates (figure 1 (Left)) but with the added constraint that $D_{\text{qat}}\leq D_{\text{total}}$. The optimal fit in our setup yields $a=6.7297$.
@@ 4.2 Loss Scaling Law
We fit the proposed equation for 757 total QAT experiments directly using Huber loss (Huber, 1964) and gradient descent optimization—a setup consistent with that of Hoffmann et al. (2022b); Chen et al. (2025b). The results are highly dependent on initialization; therefore, we select the best fit out of many random initializations. We achieve similar fit quality across different bit-widths: $R^{2}=0.982$ for 1-bit QAT and $R^{2}=0.991$ for 6-bit QAT, where $R^{2}$ is the coefficient of determination. Full fit metrics are presented in table 1. We present the fitted formula and its visualizations in appendix 12 and plot the 3D loss scaling law surface for fixed model size in figure 4. Also, in appendix 14 we fit formulas independently for each bit-width $B$ as a baseline and verify that the unified formula achieves similar fit metrics. Additional scaling law fit notes are provided in appendix 13, and appendix 18 verifies that optimal QAT fraction prediction generalizes to larger model sizes (2.2B).
@@
The fitted formulas are analytically sound: with an increase of either $D_{\text{fp}}$ or $D_{\text{qat}}$ while the other is fixed, the total loss decreases. Additionally, the proposed form effectively captures the optimal QAT fraction. From experimental results and the fitted loss function, we observe that low-bit QAT is more sensitive to the QAT fraction being optimal. Loss increase for sub-optimal QAT is higher for 1-bit than for 6-bit. Therefore, selecting the optimal QAT fraction is especially important in low-bit settings. To assess the generality of our formulation, we reproduce our findings with different pre-training and QAT hyperparameters and the SlimPajama (Soboleva et al., 2023) dataset in appendix 17.
Quantifying Sub-Optimality.
After obtaining loss scaling laws, we can analyze how the optimal QAT setup compares to a sub-optimal one. To do so, we calculate wasted token count—the number of tokens that are effectively wasted by a sub-optimal QAT fraction. Wasted token count is summarized in figure 4 for different bit widths and token counts; we use 10% QAT as the reference for the sub-optimal setup. Two factors influence the wasted token magnitude: closeness of 10% to the optimal QAT fraction and the overall flatness of the loss scaling law for high token counts. If the predicted loss is generally flat for some token count, then even high deviation from optimality will yield minor wasted token count. In the extreme case, for 1-bit QAT, the same loss can be achieved with just around 50% of compute if the optimal QAT fraction is used. This effect is still present for 2–4-bit QAT but becomes relatively small for 6-bit.
@@
Figure 3: Visualization of fitted loss scaling law for 759M model, 1-bit QAT, and different $D_{\text{qat}},D_{\text{fp}}$. Orange lines represent constant $D_{\text{total}}=D_{\text{qat}}+D_{\text{fp}}$ levels, and stars represent loss minima for each such level. It is clearly seen that the loss structure yields an optimal QAT fraction for a specific $D_{\text{total}}$. The overall phenomenon is consistent with what was discussed in section 4.1.
@@
Note on Compute Budget/Token Budget Duality.
So far, we have considered token count to be identical to compute budget as compute scales linearly with training token count. However, one may argue that since QAT employs additional operations, its complexity is higher than FP training. As QAT overhead depends only on model size, it becomes negligible with sufficiently large batch size and sequence length (appendix 19). Still, in setups where QAT overhead is substantial, one can obtain compute-based optimal QAT fraction from the token-based approach by making a substitution $D_{\text{qat}}=\frac{1}{r}\cdot D_{\text{qat}}^{\prime}$, where $r>1$ is the QAT overhead factor in the specific setup and $D_{\text{qat}}^{\prime}$ is the overhead-aware QAT token count. This will make QAT steps “more expensive” from a loss minimization perspective. In this setup, $D_{\text{fp}}+D_{\text{qat}}^{\prime}=const$ represents not iso-token levels, but rather iso-flop levels. Therefore, the inferred optimal QAT fraction will be adjusted to account for the overhead.
5 When Does QAT Restore FP Model Accuracy?
An interesting question to analyze is how QAT loss relates to FP model loss for the same total token count. Specifically, a practically significant question is “when does QAT restore FP model accuracy?” To answer this question, we plot the difference in perplexities between QAT and FP models for each total token count; appendix 13 describes how we obtain the full-precision model loss scaling law. Figure 5 presents such a plot for models of two different sizes.
@@ 5 When Does QAT Restore FP Model Accuracy?
Figure 5: Difference in perplexity between FP loss scaling law and QAT loss scaling law for two model sizes. For QAT, the loss corresponding to the optimal QAT fraction is used. Values below 0 correspond to QAT performing better than the FP model. It is clearly observed that the ability of QAT to match FP loss is greatly influenced by model size and token count. In particular, larger models are able to tolerate lower QAT precision for higher total token count budgets. Additional plot information is present in appendix 15.
6 Parameter–Precision Trade-Off
An interesting question to analyze is “for a fixed model memory requirement, how should one select QAT precision and parameter count?”. That is, to fit a specified memory constraint, one can choose high precision at the expense of lower parameter count or vice versa. This question is practically important as LLM inference is largely bottlenecked by memory bandwidth (Davies et al., 2025; Recasens et al., 2025; Dao et al., 2022). We can derive such optimality from the fitted loss scaling law.
@@ 6 Parameter–Precision Trade-Off
The results are presented in figure 6. It is clearly seen that for a fixed memory budget, optimal QAT precision decreases with training FLOP growth. This suggests that for achieving an optimal quantized model within some memory and training compute budget, one should select parameter count in advance accordingly. We believe this finding to be important for practitioners trying to achieve the best-accuracy model within memory constraints. Figure 6 is verified experimentally in appendix 16.
7 Cooldown & QAT Fusion
Section 4 revealed the importance of advance planning for QAT, accounting for the optimal fraction. This is possible only when one controls the entire pretraining process: both QAT and FP. In this context, it may be worth adjusting the training procedure to make QAT more efficient. Specifically, in this section we focus on modifications to learning rate scheduling techniques. Currently, a classic way of training models is to perform full FP training with learning rate cooldown, and then start QAT with learning rate re-warmup. We used such a setup for the scaling law in section 4 as it is universally adopted. However, a more optimal scheme may exist.
@@ 7 Cooldown & QAT Fusion
Wen et al. (2024) show that re-initializing WSD from a post-cooldown checkpoint rather than from a constant stage yields better results. However, we believe the behavior might be different when resuming training from a checkpoint with QAT. We propose a novel idea for QAT: QAT & learning rate cooldown fusion. Motivated by the idea that learning rate cooldown performs low-magnitude adjustments to weights, we speculate that a substantial part of updates during learning rate cooldown gets destroyed by QAT initialization, which, in essence, discards high-precision information. Therefore, we analyze a setup where QAT is started directly from the learning rate constant stage and learning rate cooldown is performed jointly with QAT. A schematic representation of the two schemes is presented in figure 7.
@@
Figure 7: Comparison between two different QAT schemes. In both setups, the QAT fraction is 40%. Red-shaded areas indicate zones with lowered learning rate, which we expect to correspond to minor weight updates that get effectively ignored by QAT initialization. On the left, classic QAT scheme visualization: QAT follows fully completed FP training that ends with 20% (of FP training length) learning rate decay. For QAT, the learning rate follows a cosine shape with 5% re-warmup phase. On the right, the cooldown & QAT fusion scheme is displayed. QAT starts directly from the constant learning rate stage with small re-warmup, effectively resuming the FP learning rate scheduler as if QAT was not present at all. QAT ends with 20% cooldown (of total training length). As QAT follows the classic FP learning rate recipe with usual cooldown, we call this approach cooldown & QAT fusion.
We run experiments for different model sizes and 4-bit QAT with the described cooldown & QAT fusion scheme, taking experiments with the classic QAT scheme and optimal QAT fraction as baselines. The results are displayed in table 2. In addition to perplexity, we report loss change in “wasted tokens” units. This is the total token distance between corresponding loss points in the scaling law for an optimal QAT fraction. Such a metric is reported for better impact understanding, as small perplexity differences are harder to achieve with high overall token counts. We achieve improvements across all model sizes for 4- and 6-bit widths and all token counts. Results differ for 1- and 2-bit settings; we believe this is due to the large optimal QAT fraction, which makes the FP fraction small and, consequently, the impact of cooldown & QAT fusion lower. Full results are available in appendix 20. While perplexity differences may seem small, judging the difference from the perspective of token count difference is significant. This implies substantial improvements in terms of training cost.
Table 2: Accuracy comparison between the classic QAT scheme and the cooldown & QAT fusion training scheme. The loss difference is reported in “wasted tokens”—the difference in total token count between optimal QAT fraction loss points in the loss scaling law. Substantial improvements are noticeable across different model sizes and token counts.
Perplexity Wasted tokens
Unfused Fused Unfused total tokens, %
$\bm{B}$ Model size Total tokens

### v2

@@ 1 Introduction
As Large Language Models (LLMs) grow in size and on-device applications gain traction (Xu et al., 2024), significant attention has been devoted to reducing inference costs via model compression (Frantar et al., 2022; Lin et al., 2024; Ma et al., 2023). One state-of-the-art method is quantization-aware training (QAT) (Chen et al., 2025a; Lin et al., 2024; Liu et al., 2025; Jacob et al., 2018). To adapt the model to the loss of numerical precision, QAT incorporates quantization directly into the model training process. It has been shown that QAT outperforms post-training quantization (PTQ) (Xiao et al., 2023; Banner et al., 2019), where quantization is applied after training is completed. Moreover, Liu et al. (2025) demonstrated that the best accuracy is achieved when a QAT phase follows a full-precision (FP) training phase.
@@
  • We propose a novel approach: QAT & Learning Rate Cooldown Fusion—a scheme where learning rate decay is performed jointly with quantization-aware training, eliminating redundant full-precision updates and achieving better accuracy for the same token count.
@@ 2 Related Work
Multiple works have previously addressed the problem of predicting final model loss ($L$) as a function of parameter count ($N$) and consumed tokens ($D$) (Bi et al., 2024; Hoffmann et al., 2022; Kaplan et al., 2020). The Chinchilla (Hoffmann et al., 2022) loss model is one of the most commonly used: $L(N,D)=E+AN^{-\alpha}+CD^{-\beta},$ where $A$, $C$, $\alpha$, $\beta$, and $E$ are fitted parameters. Bi et al. (2024) expand on this idea, fitting accuracy as a function of used non-embedding FLOPs (FLOP estimation of model inference without embedding layer calculations), showing that such an approach works better across different model sizes. Additionally, they show that scaling laws are greatly influenced by data quality.
@@
3 Optimal QAT Compute Allocation
To study how loss changes for different combinations of QAT/FP training length, we train models of different sizes ($N$), different FP stage token counts ($D_{\text{fp}}$), QAT token counts ($D_{\text{qat}}$), and different QAT bit widths ($B$). For the smallest model (86.0M parameters), we conduct experiments from 2.3B to 1.4T total tokens, while for the largest model (759.0M parameters), we conduct experiments from 8.5B to 669.2B total tokens. We focus primarily on 1-, 2-, 4-, and 6-bit quantization widths. Full description of our experimental setup is described in appendix A, and token counts and QAT fractions used are reported in appendix N. We also verify that the obtained post-QAT models maintain reasonable accuracy and present a comparison to full-precision models in appendix P. Specifically, our 4- and 6-bit setups achieve quality close to that of the full-precision model for the same total token count, and the drop in quality for 1- and 2-bit is reasonable.
@@ 3.1 Predicting the Optimal QAT Fraction
In this section, we focus on fitting the optimal QAT fractions directly. To account for different QAT bit widths used, we introduce the tokens-per-parameter-byte statistic. This choice was made based on several observations: larger models are generally easier to quantize, models trained for longer are harder to quantize, and smaller QAT bit widths are harder to quantize as well. While being intuitive from a QAT accuracy perspective, it can also predict the optimal fraction with high precision. Figure 2 provides a comparison between the two approaches. It is clearly seen that using tokens-per-parameter-byte provides an interpretable adjustment, facilitating a better fit.
For optimal QAT fraction prediction, we fit a function of the form
@@
where $\widehat{f}$ is the predicted optimal QAT fraction for the total tokens-per-parameter-byte count $S_{\text{total}}$, and $a$ is a fitted parameter. This function choice was made due to the observed almost linear dependency between $S_{\text{total}}$ and optimal $S_{\text{qat}}$ in log-log coordinates, but with the added constraint that $D_{\text{qat}}\leq D_{\text{total}}$. The optimal fit in our setup yields $a=6.7297$.
@@
Takeaway 1 Optimal QAT compute allocation fraction is not stationary but grows with total training tokens-per-parameter-byte ($S_{\text{total}}$) and can be predicted from it.
@@ 3.2 Loss Scaling Law
We fit the proposed equation for 757 total QAT experiments directly using Huber loss (Huber, 1964) and gradient descent optimization—a setup consistent with that of Hoffmann et al. (2022); Chen et al. (2025b). The results are highly dependent on initialization; therefore, we select the best fit out of many random initializations. We achieve similar fit quality across different bit-widths: $R^{2}=0.982$ for 1-bit QAT and $R^{2}=0.991$ for 6-bit QAT, where $R^{2}$ is the coefficient of determination. Full fit metrics are presented in table 1. We present the fitted formula and its visualizations in appendix D and plot the 3D loss scaling law surface for fixed model size in figure 4. Also, in appendix G we fit formulas independently for each bit-width $B$ as a baseline and verify that the unified formula achieves similar fit metrics. Additional scaling law fit notes are provided in appendix F, appendix K verifies that optimal QAT fraction prediction generalizes to larger model sizes (2.2B), and appendix Q conducts uncertainty analysis and parameters significance tests.
@@
The fitted formulas are analytically sound: with an increase of either $D_{\text{fp}}$ or $D_{\text{qat}}$ while the other is fixed, the total loss decreases. Additionally, the proposed form effectively captures the optimal QAT fraction. From experimental results and the fitted loss function, we observe that low-bit QAT is more sensitive to the QAT fraction being optimal. Loss increase for sub-optimal QAT is higher for 1-bit than for 6-bit. Therefore, selecting the optimal QAT fraction is especially important in low-bit settings. To assess the generality of our formulation, we reproduce our findings with different pre-training and QAT hyperparameters and the SlimPajama (Soboleva et al., 2023) dataset in appendix J.
Takeaway 2 Final loss after QAT can be accurately predicted from $N,D_{\text{qat}},D_{\text{fp}}$, and $B$ for various settings using a single formula. Moreover, the proposed loss scaling law effectively captures the phenomena of optimal QAT fraction and can be used to infer it.
4 Loss Scaling Law Implications
The unified loss scaling law obtained in the previous section allows us to analyze practically important QAT properties. In this section, we address previously unanswered questions such as: ”How bad is sub-optimal QAT compute allocation?”, ”When does QAT match FP accuracy?”, and ”How should one select QAT precision and parameter count?”
4.1 Evaluating Sub-Optimal QAT Fraction
Using the loss scaling law, we compare the optimal QAT setup with a sub-optimal one. To do this, we calculate wasted token count—the number of tokens effectively wasted by a sub-optimal QAT fraction. Figure 4 summarizes wasted token count for different bit widths and token counts; we use 10% QAT as the reference for the sub-optimal setup. Formally, with fitted loss model $L(N,D_{\text{qat}},D_{\text{fp}},B)$, we can find such $D_{\text{total}}^{*}$ and optimal $D_{\text{qat}}^{*}(N,D_{\text{total}}^{*},B)$ that achieve the same loss $l$ as sub-optimal $D_{\text{qat}}^{\text{subopt}}$ and $D_{\text{total}}^{\text{subopt}}$. Specifically:
$l$ $=L(N,D_{\text{qat}}^{\text{subopt}},D_{\text{total}}^{\text{subopt}}-D_{\text{qat}}^{\text{subopt}},B),$
$D_{\text{total}}^{*}$ $=\operatorname*{arg\,min}_{\begin{subarray}{c}D_{\text{total}}^{\prime}\in{\mathbb{N}}\\ D_{\text{qat}}^{\prime}=D_{\text{qat}}^{*}(N,D_{\text{total}}^{\prime},B)\end{subarray}}\left|L(N,D_{\text{qat}}^{\prime},D_{\text{total}}^{\prime}-D_{\text{qat}}^{\prime},B)-l\right|,$
$D_{\text{wasted}}$ $=D_{\text{total}}^{\text{subopt}}-D_{\text{total}}^{*},$
and the reported percentage in figure 4 is the fraction of total tokens: $\frac{D_{\text{wasted}}}{D_{\text{total}}^{\text{subopt}}}$.
Two factors influence the wasted tokens magnitude: closeness of 10% to the optimal QAT fraction and the overall flatness of the loss scaling law for high token counts. If the predicted loss is generally flat for some token count, then even high deviation from optimality will yield a minor wasted token count. In the extreme case, for 1-bit QAT, the same loss can be achieved with just around 50% of compute if the optimal QAT fraction is used. This effect is still present for 2–4-bit QAT but becomes relatively small for 6-bit.
@@ 4.1 Evaluating Sub-Optimal QAT Fraction
Figure 3: Visualization of the fitted loss scaling law for a 759M model, 1-bit QAT, and different $D_{\text{qat}},D_{\text{fp}}$. Orange lines represent constant $D_{\text{total}}=D_{\text{qat}}+D_{\text{fp}}$ levels, and stars represent loss minima for each such level. It is clearly seen that the loss structure yields an optimal QAT fraction for a specific $D_{\text{total}}$. The overall phenomenon is consistent with what was discussed in section 3.1.
@@
Takeaway 3 Suboptimal QAT compute allocation significantly impacts final model performance, especially in low-bit settings. In the extreme case, for 1-bit QAT, the same loss can be achieved with just around 50% of compute if the optimal QAT fraction is used.
Note on Compute Budget–Token Budget Duality.
So far, we have considered token count to be identical to compute budget as compute scales linearly with training token count. However, one may argue that since QAT employs additional operations, its complexity is higher than FP training. As QAT overhead depends only on model size, it becomes negligible with sufficiently large batch size and sequence length (appendix L). Still, in setups where QAT overhead is substantial, one can obtain compute-based optimal QAT fraction from the token-based approach by making a substitution $D_{\text{qat}}=\frac{1}{r}\cdot D_{\text{qat}}^{\prime}$, where $r>1$ is the QAT overhead factor in the specific setup and $D_{\text{qat}}^{\prime}$ is the overhead-aware QAT token count. This will make QAT steps “more expensive” from a loss minimization perspective. In this setup, $D_{\text{fp}}+D_{\text{qat}}^{\prime}=const$ represents not iso-token levels, but rather iso-FLOP levels. Therefore, the inferred optimal QAT fraction will be adjusted to account for the overhead.
4.2 When Does QAT Match FP Accuracy?
We plot the difference in perplexities between QAT and FP models for each total token count. Appendix F describes how we obtain the full-precision model loss scaling law. In summary, we incorporate full-precision training results into the fit with $B=16$, which not only allows us to predict full-precision model performance but also serves as a regularization for the fit. Figure 5 presents such a plot for models of two different sizes.
@@ 4.2 When Does QAT Match FP Accuracy?
Takeaway 4 The proposed loss scaling law effectively captures the empirical observation that larger-parameter models can tolerate lower-bit QAT. Moreover, using the loss scaling law, one can predict a range of total token counts for which QAT accuracy will not differ significantly from that of a FP model.
@@
Figure 5: Difference in perplexity between FP loss scaling law and QAT loss scaling law for two model sizes. For QAT, the loss corresponding to the optimal QAT fraction is used. Values below 0 correspond to QAT performing better than the FP model. It is clearly observed that the ability of QAT to match FP loss is greatly influenced by model size and token count. In particular, larger models are able to tolerate lower QAT precision for higher total token count budgets. Additional plot information is present in appendix H.
4.3 Parameter–Precision Trade-Off
@@
An interesting question to analyze is “for a fixed model memory requirement, how should one select QAT precision and parameter count?”. That is, to fit a specified memory constraint, one can choose high precision at the expense of a lower parameter count or vice versa. This question is practically important as LLM inference is largely bottlenecked by memory bandwidth (Davies et al., 2025; Recasens et al., 2025; Dao et al., 2022). We can derive such optimality from the fitted loss scaling law. The results are presented in figure 6. It is clearly seen that for a fixed memory budget, optimal QAT precision decreases with training FLOP growth. This suggests that for achieving an optimal quantized model within some memory and training compute budget, one should select the parameter count in advance accordingly. We believe this finding to be important for practitioners trying to achieve the best-accuracy model within memory constraints. Figure 6 is verified experimentally in appendix I.
Takeaway 5 The proposed loss scaling law can predict what QAT bit width is optimal for a fixed training compute budget and model memory footprint. It is revealed that for a fixed memory budget, optimal QAT precision decreases with training FLOP growth.
5 Beyond QAT Compute Fraction: QAT & Cooldown Fusion
Section 3 revealed the importance of advance planning for QAT, accounting for the optimal fraction. This is possible only when one controls the entire pretraining process: both QAT and FP. In this context, it may be worth adjusting the training procedure to make QAT more efficient. Specifically, in this section, we focus on modifications to learning rate scheduling techniques. Currently, a classic way of training models is to perform full FP training with learning rate cooldown, and then start QAT with learning rate re-warmup. We used such a setup for the scaling law in section 3 as it is universally adopted. However, a more optimal scheme may exist.
@@ 5 Beyond QAT Compute Fraction: QAT & Cooldown Fusion
Wen et al. (2024) show that re-initializing WSD from a post-cooldown checkpoint rather than from a constant stage yields better results. However, we believe the behavior might be different when resuming training from a checkpoint with QAT. We propose a novel idea: QAT & Learning Rate Cooldown Fusion. Motivated by the idea that learning rate cooldown performs low-magnitude adjustments to weights, we speculate that a substantial part of updates during learning rate cooldown gets destroyed by QAT initialization, which, in essence, discards high-precision information. Therefore, we analyze a setup where QAT is started directly from the learning rate constant stage and learning rate cooldown is performed jointly with QAT. A schematic representation of the two schemes is presented in figure 7.
@@
Figure 7: Comparison between two different QAT schemes. In both setups, the QAT fraction is 40%. Red-shaded areas indicate zones with lowered learning rate, which we expect to correspond to minor weight updates that get effectively ignored by QAT initialization. On the left, classic QAT scheme visualization: QAT follows fully completed FP training that ends with 20% (of FP training length) learning rate decay. For QAT, the learning rate follows a cosine shape with 5% re-warmup phase. On the right, the ”QAT & Learning Rate Cooldown Fusion” scheme is displayed. QAT starts directly from the constant learning rate stage with small re-warmup, effectively resuming the FP learning rate scheduler as if QAT was not present at all. QAT ends with 20% cooldown (of total training length). As QAT follows the classic FP learning rate recipe with usual cooldown, we call this approach QAT & Learning Rate Cooldown Fusion.
We ran experiments with different model sizes and 4-bit QAT using the described ”QAT & Cooldown Fusion” scheme, taking experiments with the classic QAT scheme and optimal QAT fraction as baselines. The results are shown in table 2. In addition to perplexity, we report loss change in “wasted tokens” units. This is the total token distance between corresponding loss points in the scaling law for an optimal QAT fraction. Such a metric is reported for better impact understanding, as small perplexity differences are harder to achieve with high overall token counts. We achieved improvements across all model sizes for 4- and 6-bit widths and all token counts. Results differed for 1- and 2-bit settings; we believe this is due to the large optimal QAT fraction, which makes the FP fraction small and, consequently, the impact of QAT & cooldown fusion lower. Full results are available in appendix M. While perplexity differences may seem small, judging the difference from the perspective of token count difference is significant. This implies substantial improvements in terms of training cost.
Table 2: Accuracy comparison between the classic QAT scheme and the ”QAT & Learning Rate Cooldown Fusion” training scheme. The loss difference is reported in “wasted tokens”—the difference in total token count between optimal QAT fraction loss points in the loss scaling law (formally defined in appendix O). Substantial improvements are noticeable across different model sizes and token counts.
[…]
Unfused (baseline) Fused (ours) Unfused total tokens, %
$\bm{B}$ Model size, M $\bm{D_{\text{total}}}$
@@
Takeaway 6 The proposed method of learning rate scheduling, ”QAT & Learning Rate Cooldown Fusion,” further improves QAT efficiency by adjusting the training procedure beyond changing the QAT fraction. While being practically useful, this method also suggests that modifications to the universally accepted QAT pipeline can further improve QAT efficiency.